OpenStudy (hopeless):

David is twice as old as his daughter. In 14yr., David's age will be three times what his daughter's age was 6 yrs. ago. How old are they now?

OpenStudy (anonymous):

David = x, Daughter = y x = 2y x+14 = 3(y-6) = 3y - 18 x = 3y - 32 -x = -2y 0 = y - 32 y = 32 x = 64